OpenStudy (anonymous):

Use the fOIL method to find the product of (7x^8-4)(x^9-5)

OpenStudy (anonymous):

\[(7x^8-4)(x^9-5)\] \[=(7x^8\times x^9)+(7x^8 \times -5)+(-4\times x^9) +(-4\times -5)\]\[=7x^\left( 17 \right)-35x^8-4x^9+20\]\

OpenStudy (anonymous):

you know how to use foil right...and this is algerbra because when u mutltiply variable's by one another you add the terms fro example the x^8*x^9=x^8+9=x^17

OpenStudy (anonymous):

and to use foil u multiply the first term in the frist bracket by both terms in the 2nd bracket and you do the same with the second term in the first bracket just stating the facts :) hope this helps